OpenStudy (anonymous):

(14x^3 - 21x^2 + 28x) / (7x) Help me?

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@mossyfish

OpenStudy (anonymous):

Hint:\[14 x^3-21 x^2+28 x=7 x \left(2 x^2-3 x+4\right) \]

OpenStudy (anonymous):

So, let's start off by dividing everything by 7. Forget about the variables for now. 14/7=? 21/7=? 28/7=?

OpenStudy (anonymous):

is it 21x^4 - 28x^3 + 35x^2??

OpenStudy (anonymous):

We're dividing :)

OpenStudy (anonymous):

Well, 14/7 = 7 21/7 = 3 28/7 =4

OpenStudy (anonymous):

14/7 is not 7. We can check this by multiplying 7 by 7, we get 49. You're right that 21/7=3 and that 28/7 is 4 :)

OpenStudy (anonymous):

Sorry, it's 2. I was thinking subtraction.

OpenStudy (anonymous):

That's alright!! So, now that we've divided the 7 part, let's divide by x :) What is 2x^3/x 3x^2/x 4x/x ?

OpenStudy (anonymous):

I don't know how to do that...

OpenStudy (anonymous):

Perfectly okay :) When we divide a variable by the same variable, we just take it away. I'll do the first one as an example. 2x^3 means we have x*x*x. If we divide by 1 x, we now have x*x. It's 2x^2 Does this make sense? This website goes a little more in depth about dividing variables http://www.dummies.com/how-to/content/dividing-variables-in-algebra.html

OpenStudy (anonymous):

So the second one would be 3x^1?

OpenStudy (anonymous):

Exactly! Except, we don't need to have that ^1 there because it's indicating that there's only one x... So it's just 3x :)

OpenStudy (anonymous):

So the last one would be just 4x too?

OpenStudy (anonymous):

The last one has one x to start with, so it'd be just 4 :)

OpenStudy (anonymous):

So, great, we have all of our numbers divided! Let's plug this back into the equation. \[2x^2-3x+4\]

OpenStudy (anonymous):

Would that be it? or do we have to solve it from there?

OpenStudy (anonymous):

Unless if we have a value for x, we can't really solve any further. So, that'd be it!! Great job